A golf ball is hit so that when it is directly above a point that is x feet from the golfer, it is
h(x)= 30x-1/10x^2 feet above the ground
How far from the golfer will the ball hit the
ground?
(A) 100 feet (B) 150 feet (C) 200 feet (D) 250 feet (E) 300 feet

Answer by lwsshak3(11628)

A golf ball is hit so that when it is directly above a point that is x feet from the golfer, it is
h(x)= 30x-1/10x^2 feet above the ground
How far from the golfer will the ball hit the
ground?
**
when the ball hits the ground, h(x)=0
30x-1/10x^2=0
-0.1x^2+30x=0
-x^2+300x=0
-x(x-300)=0
-x=0 (reject)
x-300=0
x=300
How far from the golfer will the ball hit the
ground? 300 ft, ans (E)